The sum of two numbers is 2. & product of the numbers is 3 find the sum of thier reciprocals

Grade:9

2 Answers

Arun
6 years ago
Let the numbers be a and b
Then 
a+ b = 2        (i)
a*b = 3          (ii)
Now we are asked 
1/a + 1/b. =?
Now 
If we divide equation i by equation ii
Then we get 
1/a + 1/b  = 2/3
This is the answer
Harshit Panwar
6 years ago
x+y=2
x × y=3
1/x+1/y
=(x+y)/x × y
=2/3
=0.6666.......
2/3 or 0.6666.... Is the answer.
Hope it helps you....